Complete the ordered pair (4) so that it is a solution of x + 2y = 9
Answer by stanbon(75887)

Complete the ordered pair (4) so that it is a solution of x + 2y = 9
-----
If x = 4, 4 + 2y = 9
2y = 5
y = (5/2)
----------------
Ans: (4,(5/2))
=====================
Cheers,
Stan H.